A block weight 200N rest on a rough horizontal table. When the block just begins to slide the pulling force is found to be 40N. What is the coefficient of friction between the two surfaces
1
Expert's answer
2019-07-09T11:21:51-0400

Applying the Newton’s Second Law of Motion, we get (at the moment when the block begins to slide the acceleration is zero):


"\sum F_x = ma_x = 0,""F_{pull} - F_{s.fr} = 0,""F_{pull} = F_{s.fr} = \mu_s N = \mu_s W,"

here, "F_{pull} = 40N" is the pulling force required to start the block of weight "W = 200 N" to slide, "F_{s.fr}" is the force of static friction, "\mu_s" is the coefficient of static friction, "N" is the normal force.

Then, from the last formula we can calculate the coefficient of static friction:


"\mu_s = \dfrac{F_{pull}}{W} = \dfrac{40 N}{200 N} = 0.2"

Answer:

"\mu_s = 0.2"